OpenStudy (anonymous):

PLEASE HELPPP!!! 1. Choose the equation that could be used to solve the given number problem. The sum of two consecutive integers is 141. Find the integers. (Points : 1) n + 1= 141 2n + 2 = 141 2n = 141 2n + 1 = 141 2. Choose the equation that could be used to solve the given number problem. The sum of two consecutive odd integers is 112. Find the integers. (Points : 1) 2n + 3 = 112 2n + 2 = 112 2n + 1 = 112 n + 2 = 112

OpenStudy (anonymous):

For first question : If sum of two numbers was Even one of the numbers wasnt Even . Look : 2+1=3 + e = e Got it ?

OpenStudy (anonymous):

is that the formula?

OpenStudy (anonymous):

So if a number wanst even * in 2 so to find the number isnt even 2n use :) OK ?

OpenStudy (anonymous):

1. two consecutive integers...n and n + 1 n + n+ 1 = 141 can you take it from here by combining like terms ?

OpenStudy (anonymous):

2. two consecutive odd integers...n and n + 2 now just do it like the first problem

OpenStudy (anonymous):

Is the first one C and the second one B? @kelliegirl33 @E.ali

OpenStudy (anonymous):

or are both B?

OpenStudy (anonymous):

No !

OpenStudy (anonymous):

first one is not b or c.

OpenStudy (anonymous):

second one is b

OpenStudy (anonymous):

So the first one is D

OpenStudy (anonymous):

yep !

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

thank youuuu!!!!!

OpenStudy (anonymous):

your welcome ! Thank u @kelliegirl33!

OpenStudy (anonymous):

you deserved it :)

OpenStudy (anonymous):

THANKS !